OpenStudy (gracie):

how do u do that 7c3d2/21c2d3

OpenStudy (katrinakaif):

I have replied to this on the other question. Please Refer =)

myininaya (myininaya):

7/21=1/3 c^3/c^2=c d^2/d^3=1/d